OpenStudy (anonymous):

PLEASE HELP! Write an equation of a line in slope intercept form that is parallel to y = 3x+6 and passes through the point (-10, 2.5)

OpenStudy (anonymous):

y-y=m(x-x) Plug in your values of x and y. y-2.5=m(x+10) Take the slope of y=3x+6 and plug it in where the m is. y-2.5=3(x+10)